# size of isogeny classes of abelian varieties in the extended lubin-tate case

### v1

@@ 4.2. Local conditions away from $p$
The above proposition proves that, if $R$ is an order that occurs as an endomorphism ring of an abelian variety isogenous to $A$, then $R$, must be maximal at $p.$ Consequently, it differs from the maximal order or more generally, any order that is maximal at $p$ at only finitely many primes away from $p.$ We know that since $R$ contains the Frobenius endomorphism, hence the local order ${\mathbb{Z}}_{l}[\alpha]\subset R_{l}$. Therefore, in order for $R$ to have the “smallest” local conditions, we will require ${\mathbb{Z}}_{l}[\alpha]=R_{l}.$ The following Lemma proves that we can always adjust the local conditions of an order at $l\neq p$ up to isogeny. Thus, it guarantees the existence of an abelian variety in the isogeny class over ${\mathbb{F}}_{q}$ with endomorphism ring $R$ with the local conditions specified at primes other than $p$.
@@
Let $R$ be an order in $K$ containing ${\mathbb{Z}}[\alpha].$ We fix an endomorphism algebra $K$ for an isogeny class of abelian varieties. Let $A$ over ${\mathbb{F}}_{q}$ be an element of the isogeny class such that $\operatorname{End}(A)=R^{\prime}$. Then we can find an abelian variety $A^{\prime}$ isogenous to $A$ over ${\mathbb{F}}_{q}$ such that $\operatorname{End}(A^{\prime})\otimes{\mathbb{Z}}_{l}=R_{l}$ for all $l\neq p.$
The proof is already given in [Wat69]. But for the sake completeness we present it here as well. We denote by $T_{l}(A)$ the Tate module of $A$ and let $V_{l}(A)=T_{l}(A)\otimes{\mathbb{Q}}_{l}.$
@@
Without loss of any generality we assume $R^{\prime}\subset R$. For primes larger than the index of $R^{\prime}$ in $R$, the two orders are locally the same. Let $l_{1},\dots,l_{m}$ be finitely many primes where the two orders are not the same. According to Tate’s theorem $R^{\prime}\otimes{\mathbb{Z}}_{l}$ takes the lattice $T_{l}(A)$ to $T_{l}(A)$ in $V_{l}(A)$ for all primes $l.$ Since $K$ is commutative, $V_{l}(A)$ is a rank one module over $K\otimes{\mathbb{Q}}_{l}.$ Therefore, for any order in $K$, there exists a lattice with that prescribed order as its endomorphism ring. Let $T_{l_{1}}$ be a sub-lattice of $T_{l}(A)$ that is preserved by the order $R\otimes{\mathbb{Z}}_{l_{1}}$. The quotient of the two lattices corresponds to a finite index subgroup of $A.$ We quotient out $A$ by it to get an abelian variety $A^{\prime}$ isogenous to $A$ such that $T_{l_{1}}(A^{\prime})=T_{l_{1}}(A).$ Now repeat this process for all the other primes to finish the proof. ∎
Remark 4.1.
We would like to clarify that from the above Lemma, it would seem that $\operatorname{End}A$ only depends on its localisation at the primes. However, as Waterhouse remarks in [Wat69], the isomorphism type of the order is not determined by the local conditions. That is, we could have two or more orders that are isomorphic locally but may not be isomorphic as orders in $K.$
We summarise the local conditions in the following notation that characterises the orders in endomorphism ring of the isogeny class with “smallest possible local conditions”.
Notation 4.3.
We define $R_{n}\subset K$ to be the order satisfying the following conditions.
@@
    It is maximal at $p,$ that is, $R_{n}\otimes{\mathbb{Z}}_{p}={\mathcal{O}}_{K_{1}}\oplus{\mathcal{O}}_{K_{2}}$
@@
    At the prime $l,$ we have $R_{n}\otimes{\mathbb{Z}}_{l}={\mathbb{Z}}_{l}[\alpha].$
5. Properties of abelian varieties with endomorphism ring $R_{n}$.
In this section, we note certain useful properties of abelian varieties with endomorphism by $R_{n}$. The properties that we prove below are entirely determined by the endomorphism ring $R_{n}$ and will be useful for us in subsequent sections.
@@ 5.1. Field of definition
(First property) Let $B$ be an abelian variety over $k$. Let $R_{n}$ be its the endomorphism ring. Then the abelian variety $B$ is defined over ${\mathbb{F}}_{q^{n}}$.
@@
Without any loss of generality, we prove this result for $n=1.$ We write $R_{1}=R.$ We can assume that $B$ is isogenous to an abelian variety $B^{\prime}$ that has endomoprhism by the maximal order so that $R\subset\operatorname{End}B^{\prime}$. Let $\varphi:B^{\prime}\rightarrow B$ be such an isogeny. Therefore the kernel of $\varphi$ is invariant under $R$, in particular it is invariant under the action of $\alpha$, the $q$-Frobenius of $B$. This implies that kernel is also Galois invariant under $\operatorname{Gal}(k/{\mathbb{F}}_{q})$ because it acts by the Frobenius of $B$. This implies that $B$ is in fact defined over ${\mathbb{F}}_{q}.$
@@ 5.2. Central leaf
In this sub-section, we will see that the property of $R_{n}$ being maximal at $p$ determines the isomorphism type of the $p$-divisible group of the abelian variety over $k.$ In order to show this, we first state some definitions and results we need.
Notation 5.2.
Let $R_{k}$ denote the ring $W(k)[F,V]$. Under the contravariant Dieudonné theory, we write ${\mathscr{G}}_{n,m}$ to be the $p$-divisible group over $k$ corresponding to the Dieudonné module $R_{k}/R_{k}(F^{n}-V^{m})$, which we denote by ${\mathbb{D}}(G_{m,n})$
We note that under this characterisation, the $p$-divisible group ${\mathscr{G}}_{m,n}$ will have dimension $n$ and height $m+n.$ The following result is well known, but for the sake of completeness, we give a proof here.
@@
There exists a unique $p$-divisible group of dimension one (and co-dimension one) over $k.$
@@
Let ${\mathscr{G}}$ be a $p$-divisible group of dimension one and height $n$ over $k$. let ${\mathbb{D}}$ be its Dieudonné module. We prove the proposition for dimension one case. The case of co-dimension will follow from the uniqueness of ${\mathscr{G}}$ and the fact that ${\mathscr{G}}_{1,n}^{\vee}$, the dual ${\mathscr{G}}_{1,n}$ is ${\mathscr{G}}_{n,1}$. We know that ${\mathscr{G}}$ is isogenous to ${\mathscr{G}}_{1,n}$ over $k$. We need to show that in fact, they are isomorphic. We first note that ${\mathbb{D}}$ and ${\mathbb{D}}({\mathscr{G}}_{1,n})$ have the same rank, that is they are both of rank $n+1$ as $W(k)$-modules. We only need to show that they are isomorphic as $R_{k}$-modules. We show this by constructing a suitable $R_{k}$ bases for ${\mathbb{D}}$. Note that ${\mathbb{D}}/F{\mathbb{D}}$ is a one dimensional vector space over $k$, and therefore, spanned by (the image of) an element, say $e_{1}$ in ${\mathbb{D}}.$ Now, let $Fe_{1}=e_{2}.$ We note that $Ve_{2}=pe_{1}.$ Next, we let $Fe_{2}=e_{3}$. Note that $e_{3}$ cannot be in the span of $e_{2}$, otherwise $e_{2}$ would be in the span of $e_{1}.$ Moreover, $e_{3}$ cannot be the $W(k)$-span of $e_{1}$ and $e_{2}.$ For if $e_{3}=Fe_{2}=ae_{1}+be_{2}=ae_{1}+bFe_{1}$, then $ae_{1}=0$ in ${\mathbb{D}}/F{\mathbb{D}}$. This implies that $a=0$ which is not possible. We carry on the same process until we hit a bases such that $F$ acts as $Fe_{i}=e_{i+1}$ for $1\leq i\leq n$. Finally writing $Fe_{n+1}=a_{1}e_{1}+a_{2}e_{2}+\dots+a_{n+1}e_{n+1}$ and applying $V$ on both sides implies that $pe_{n+1}=a_{1}Ve_{1}$ and $a_{i}=0$ for $i\geq 2$. Since $e_{1}$ is not in the image of $F,$ we see that $a_{1}=p$ and hence $Fe_{n+1}=pe_{1}$. This gives us an isomorphism between ${\mathscr{G}}$ and ${\mathscr{G}}_{1,n}$.
@@
(Central leaves) Let ${\mathscr{G}}$ be a $p$-divisible group over ${\mathbb{F}}_{q}.$ Then the central leaf of ${\mathscr{G}}$ in $\mathcal{A}_{g,1}\otimes{\mathbb{F}}_{q}$ is defined as
$c({\mathscr{G}})=\{B\in\mathcal{A}_{g,1}({\mathbb{F}}_{q^{n}})\mid{\mathscr{G}}\otimes k\simeq B[p^{\infty}]\otimes k\}.$
@@
(Second property) Let $B$ be an abelian variety over ${\mathbb{F}}_{q^{n}}$ with endomorphism ring $R_{n}\subset K$. Let $c$ be the central leaf of ${\mathscr{G}}_{1,g}\times{\mathscr{G}}_{g-1,g}$. Then $B$ lies in the central leaf $c.$
@@
We know that $\operatorname{End}(B[p^{\infty}])\otimes k={\mathcal{O}}_{K_{1}}\oplus{\mathcal{O}}_{K_{2}}$. This implies that its Dieudonné module ${\mathbb{D}}\otimes k$ splits into a direct sum ${\mathbb{D}}_{1}\oplus{\mathbb{D}}_{2}$ such that $\operatorname{End}{\mathbb{D}}_{1}$ and $\operatorname{End}{\mathbb{D}}_{2}$ are maximal at $p.$ This gives the desired decomposition of $B[p^{\infty}]\otimes k$ into a product ${\mathscr{G}}_{1}\times{\mathscr{G}}_{2}$. We know that ${\mathscr{G}}_{1}\times{\mathscr{G}}_{2}$ is isogenous to ${\mathscr{G}}_{1,g}\times{\mathscr{G}}_{g-1,g}$. But since the endomorphism ring preserves the kernel, it must also factor into a direct summand. We see that ${\mathscr{G}}_{1}$ (respectively ${\mathscr{G}}_{2}$) is isogenous, hence isomorphic to ${\mathscr{G}}_{1,g}$ (respectively ${\mathscr{G}}_{g-1,g}$). This shows that $B$ lies in the central leaf $c.$
@@ 6. Isogeny classes and the class group
Let $R_{n}\subset K$ be the order defined in Notation 4.3. Let $I\subset K$ be an invertible $R_{n}$ module, that is, $I$ is a fractional ideals of $R_{n}$ such that there exists some fractional ideal $J$ with $IJ=R_{n}$. We say two fractional ideals, $I_{1}$ and $I_{2}$ are equivalent if and only if there exists some $a\in K^{\times}$ such that $I_{1}=aI_{2}.$
@@
(Class group of an order) Under the equivalence mentioned above, we define the equivalence classes of fractional ideals of $R_{n}$ to be the class group of $R_{n}$, denoted as $cl(R_{n}).$
The heart of rest of the paper lies in the following proposition that gives us a convenient characterisation of (isogenous) abelian varieties with endomorphism ring $R_{n}$ over $L^{\prime}.$
@@
Let $\Phi$ be the primitive CM type on $K$ corresponding to $\widetilde{A}.$ Then there exists an injective map from the class group of $R_{n}$ to the set of isomorphism classes of abelian varieties over $L^{\prime}$ with endomorphism ring $R_{n}$ inside $(K,\Phi).$
@@
Let $I$ be an element of $\operatorname{cl}(R_{n}).$ We recall that $\Phi$ gives is an isomorphism $K\otimes{\mathbb{C}}={\mathbb{C}}^{\Phi}$ where ${\mathbb{C}}^{\Phi}$ denotes $g$ copies of ${\mathbb{C}}$ indexed by the embeddings in $\Phi.$ Since $I$ is a lattice in $K,$ we can embed it inside ${\mathbb{C}}^{\Phi}$. We refer to this embedding as $\Phi(I).$ The complex torus ${\mathbb{C}}^{\Phi}/\Phi(I)$ admits a polarisation (see sub-section 8) and therefore we get an abelian variety over ${\mathbb{C}}$ which we denote by $A_{\Phi(I)}$. Since $\Phi$ is primitive, $A_{\Phi(I)}$ is simple and so its endomorphism algebra is $K$. By construction $\operatorname{End}(A_{\Phi(I)})=R_{n}.$ Furthermore, since $A_{\Phi(I)}$ is a CM abelian variety, it descends, with its entire endomorphism algebra to a number field. Now, if the abelian variety had CM by the maximal order, then the abelian variety is defined over the maximal unramified extension of the reflex field $K^{*}$ and hence it would be defined over an unramified extension of ${\mathbb{Q}}_{p}.$ In this case, however, the field of definition might be slightly larger than that and might possibly be a ramified extension of ${\mathbb{Q}}_{p}$. We first rule out this possibility for the abelian variety $A_{\Phi(R_{n})}.$ We note that $A_{\Phi(R_{n})}$ is isogenous to the abelian variety $A^{\prime}$ corresponding to the lattice ${\mathcal{O}}_{K}$ in $K$. Now, $A^{\prime}$ has CM by a maximal order over ¯$\overline{{\mathbb{Q}}}_{p}.$ Let $F$ be the unramified extension of ${\mathbb{Q}}_{p}$ where $A^{\prime}$ is defined. Since the kernel of the isogeny has order equal to the index of $R_{n}$ in ${\mathcal{O}}_{K}$, we can embed it inside $A^{\prime}[m]$ for some $(m,p)=1$. Now the kernel is defined over some extension $F^{\prime}$ of $F$. This must be unramified because the inertia group $I(\overline{v})$ acts trivially on the $m$-torsion and therefore on the subgroup as well. Here $\bar{v}$ refers to the extension of a place $v$ of $F^{\prime}$ to its separable closure. This descends $A_{\Phi(R_{n})}$ to an unramified extension of ${\mathbb{Q}}_{p}.$ This same argument works in descending an abelian variety corresponding to a fractional ideal $I$ whose index in $R_{n}$ is prime to $p.$ So suppose that $P$ is a fractional ideal of $R_{n}$ with index, a power of $p$. We extend this to a fractional ideal $P^{\prime}$ of ${\mathcal{O}}_{K}$. Since $R_{n}$ is maximal at $p$, the index of $P$ in $R_{n}$ equals the index of $P^{\prime}$ in ${\mathcal{O}}_{K}$. Hence the index of $P$ in $P^{\prime}$ is prime to $p.$ Since the abelian variety corresponding to $P^{\prime}$ has endomorphism by the maximal order, it descends to an unramified extension of ${\mathbb{Q}}_{p}.$ We can now repeat above argument with $A_{\Phi(P)}$. Finally, under the equivalence of ideals in $cl(R_{n})$, two equivalent ideals would correspond to isomorphic lattices and hence give us isomorphic abelian varieties. ∎
@@
We note that all the abelian varieties constructed in 6.2 are isogenous to $\widetilde{A}_{L^{\prime}}$ over $L^{\prime}$ Lemma 3.2. Moreover, their special fibres are in fact defined over ${\mathbb{F}}_{q^{n}}$ by Lemma 5.1 and have $\alpha^{n}\in K$ as their Frobenius morphism.

### v2

@@ 3.2. Local conditions away from $p$
The above proposition proves that if $R$ is an order that occurs as an endomorphism ring of an abelian variety isogenous to $A$, then $R$ must be maximal at $p.$ Hence, it differs from the maximal order or more generally, any order that is maximal at $p$ at only finitely many primes away from $p.$ Let $l\neq p$ be any arbitrary prime. We know that since $R$ contains the Frobenius endomorphism, hence the local order ${\mathbb{Z}}_{l}[\alpha]\subseteq R\otimes{\mathbb{Z}}_{l}$ where $\alpha$ is the Weil-$q$ number of $A.$ In fact we have the following Lemma that proves that we can always adjust the local conditions of an order at $l\neq p$ up to isogeny. Thus, it guarantees the existence of an abelian variety in the isogeny class over ${\mathbb{F}}_{q}$ with endomorphism ring $R$ with the local conditions specified at primes other than $p$.
@@
Let $A$ be an abelian variety over ${\mathbb{F}}_{q}$ with endomoprhism algebra $K={\mathbb{Q}}(\alpha)$ as above. Let $R$ be an order in $K$ containing ${\mathbb{Z}}[\alpha].$ Then we can find an abelian variety $A^{\prime}$ isogenous to $A$ over ${\mathbb{F}}_{q}$ such that $\operatorname{End}(A^{\prime})\otimes{\mathbb{Z}}_{l}=R\otimes{\mathbb{Z}}_{l}$ for all primes $l\neq p.$
The proof is already given in [Wat69], Porism 4.3. For the sake completeness we present it here as well. We denote by $T_{l}(A)$ the Tate module of $A$ and let $V_{l}(A)=T_{l}(A)\otimes{\mathbb{Q}}_{l}.$
@@
Let $\operatorname{End}(A)=S$. Without loss of any generality we assume $S\subseteq R$. By Proposition 3.1, we know that $R\otimes{\mathbb{Z}}_{p}=S\otimes{\mathbb{Z}}_{p}={\mathcal{O}}_{K^{\prime}}\oplus{\mathcal{O}}_{K^{\prime}}$, where we recall the notation of $K^{\prime}$ given in the beginning of the section. Hence the index of $S$ in $R$ is not divisible by $p$. Thus it is enough to consider primes $l\neq p$. For primes larger than the index of $S$ in $R$, the two orders are locally the same. Let $l_{1},\dots,l_{m}$ be finitely many primes where the two orders are not the same. According to Tate’s theorem $S\otimes{\mathbb{Z}}_{l}$ takes the lattice $T_{l}(A)$ to $T_{l}(A)$ in $V_{l}(A)$ for all primes $l.$ Since $K$ is commutative, $V_{l}(A)$ is a rank one module over $K\otimes{\mathbb{Q}}_{l}.$ Therefore, for any order in $K$, there exists a lattice with that prescribed order as its endomorphism ring. Let $T_{l_{1}}$ be a sub-lattice of $T_{l}(A)$ that is preserved by the order $R\otimes{\mathbb{Z}}_{l_{1}}$. The quotient of the two lattices corresponds to a finite index subgroup of $A.$ We quotient out $A$ by it to get an abelian variety $A^{\prime}$ isogenous to $A$ such that $T_{l_{1}}(A^{\prime})=T_{l_{1}}(A).$ Now repeat this process for all the other primes to finish the proof. ∎
Now for each positive integer $n$, in order get a lower bound on $I(A,{\mathbb{F}}_{q^{n}})$, we count the isomorphism classes of abelian varieties in $I(A,{\mathbb{F}}_{q^{n}})$ with endomorphism ring $R_{n}$ that is characterized by the local conditions summarised below. We note that the order $R_{n}$ is the “smallest possible” at the primes $l\neq p$ and satisfies the necessary condition at the prime $p$ as proved in Proposition 3.1.
Definition 3.3.
Let $n$ be a positive integer. We define $R_{n}\subset K={\mathbb{Q}}(\alpha)$ to be the order satisfying the following conditions:
@@
    It is maximal at $p,$ that is, $R_{n}\otimes{\mathbb{Z}}_{p}={\mathcal{O}}_{K^{\prime}}\oplus{\mathcal{O}}_{K^{\prime}}$
@@
    At the prime $l,$ we have $R_{n}\otimes{\mathbb{Z}}_{l}={\mathbb{Z}}_{l}[\alpha^{n}].$
4. Properties of abelian varieties of Lubin-Tate type.
In this section, we note certain useful properties of abelian varieties with endomorphism ring characterised in Definition 3.3 that will be useful for us in subsequent sections.
@@ 4.1. Field of definition
(First property) Let $B$ be an abelian variety over $k$. Let $n$ be a positive integer and let $R_{n}$ be its the endomorphism ring as defined in Definition 3.3. Then the abelian variety $B$ is defined over the field extension ${\mathbb{F}}_{q^{n}}$ of ${\mathbb{F}}_{q}$.
@@
Without any loss of generality, we prove this result for $n=1.$ We write $R_{1}=R.$ We can assume that $B$ is isogenous to an abelian variety $B^{\prime}$ that has endomoprhism by the maximal order so that $R\subset\operatorname{End}B^{\prime}$. Let $\varphi:B^{\prime}\rightarrow B$ be such an isogeny. Therefore the kernel of $\varphi$ is invariant under $R$, in particular it is invariant under the action of $\alpha_{B}$, the $q$-Frobenius of $B$. This implies that kernel is also Galois invariant under $\operatorname{Gal}(k/{\mathbb{F}}_{q})$ because it acts by the Frobenius of $B$. This implies that $B$ is in fact defined over ${\mathbb{F}}_{q}.$
@@ 4.2. Central leaf
Let $D_{k}$ denote the ring $W(k)[F,V]$ and $\mathbb{D}(-)$ denote the contravariant Dieudonné functor. For any positive integers $a$ and $b$, we write ${\mathscr{G}}_{a,b}$ to be the $p$-divisible group over $k$ corresponding to the Dieudonné module $\mathbb{D}({\mathscr{G}}_{a,b}):=D_{k}/D_{k}(F^{a}-V^{b})$, under the Dieudonné theory. We note that under this characterisation, the $p$-divisible group ${\mathscr{G}}_{a,b}$ will have dimension $a$ and height $a+b.$ The Dieudonné-Manin classification tells us that every $p$-divisible group over $k$ is isogenous to a product of ${\mathscr{G}}_{a,b}$ as $a$ and $b$ vary over positive integers. In fact, in dimension one we have the following stronger uniqueness result.
@@
For any positive integer $h$, there exists a unique $p$-divisible group of dimension one and height $h$ over $k$. Consequently, by considering its dual the same is true for a $p$-divisible group of co-dimension one and height $h$.
@@
This result is well known. Let ${\mathscr{G}}$ be a $p$-divisible group of dimension one and height $h$ over $k$. let ${\mathbb{D}}({\mathscr{G}})$ be its Dieudonné module. Using the Dieudonné-Manin classification, we know that ${\mathscr{G}}$ is isogenous to ${\mathscr{G}}_{1,h-1}$ over $k$. We need to show that in fact, they are isomorphic. We first note that ${\mathbb{D}}({\mathscr{G}})$ and ${\mathbb{D}}({\mathscr{G}}_{1,h-1})$ have rank $h$ as $W(k)$-modules. We only need to show that they are isomorphic as $D_{k}$-modules. We show this by constructing a suitable $D_{k}$ bases for ${\mathbb{D}}({\mathscr{G}})$. Note that ${\mathbb{D}}({\mathscr{G}})/F({\mathbb{D}}({\mathscr{G}}))$ is a one dimensional vector space over $k$ and therefore, spanned by (the image of) an element, say $e_{1}$ in ${\mathbb{D}}({\mathscr{G}}).$ Now, let $Fe_{1}=e_{2}.$ We note that $Ve_{2}=pe_{1}.$ Next, we let $Fe_{2}=e_{3}$. Note that $e_{3}$ cannot be in the span of $e_{2}$, otherwise $e_{2}$ would be in the span of $e_{1}.$ Moreover, $e_{3}$ cannot lie in the $W(k)$-span of $e_{1}$ and $e_{2}.$ For if $e_{3}=Fe_{2}=ae_{1}+be_{2}=ae_{1}+bFe_{1}$, then $ae_{1}=0$ in ${\mathbb{D}}({\mathscr{G}})/F({\mathbb{D}}({\mathscr{G}}))$. This implies that $a=0$ which is not possible. We carry on the same process until we hit a bases such that $F$ acts as $Fe_{i}=e_{i+1}$ for $11\leq i\leq h-$ 1. Finally writing $Fe_{h}=a_{1}e_{1}+a_{2}e_{2}+\dots+a_{h}e_{h}$ and applying $V$ on both sides implies that $pe_{h}=a_{1}Ve_{1}$ and $a_{i}=0$ for $i\geq 2$. Since $e_{1}$ is not in the image of $F,$ we see that $a_{1}=p$ and hence $Fe_{h}=pe_{1}$. This gives us an isomorphism between ${\mathscr{G}}$ and ${\mathscr{G}}_{1,h-1}$.
@@
(Central leaves) Let ${\mathscr{G}}$ be a $p$-divisible group over ${\mathbb{F}}_{q}.$ Then the ${\mathbb{F}}_{q^{n}}$ points of the central leaf of ${\mathscr{G}}$ in $\mathcal{A}_{g,1}\otimes{\mathbb{F}}_{q}$ is defined as
$c({\mathscr{G}})({\mathbb{F}}_{q^{n}})=\{B\in\mathcal{A}_{g,1}({\mathbb{F}}_{q^{n}})\mid{\mathscr{G}}\otimes k\simeq B[p^{\infty}]\otimes k\}.$
@@
We note that $c(A)$ in fact cuts out a closed sub-variety in the Newton stratum of $A$.
@@
(Second property) Let $n$ be a positive integer. Let $B$ be an abelian variety over ${\mathbb{F}}_{q^{n}}$ satisfying the conditions mentioned at the beginning of Section 3. Let $\operatorname{End}(B)=R_{n}\subset K$. Then $B$ lies in the central leaf of the $p$-divisible group ${\mathscr{G}}_{1,g-1}\times{\mathscr{G}}_{g-1,1}$.
@@
By applying Proposition 3.1 to the abelian variety $B$, we see that $\operatorname{End}(B[p^{\infty}])\otimes k={\mathcal{O}}_{K^{\prime}}\oplus{\mathcal{O}}_{K^{\prime}}$. This implies that its Dieudonné module ${\mathbb{D}}\otimes k$ splits into a direct sum ${\mathbb{D}}_{1}\oplus{\mathbb{D}}_{2}$ such that $\operatorname{End}{\mathbb{D}}_{1}$ and $\operatorname{End}{\mathbb{D}}_{2}$ are maximal at $p.$ This gives the desired decomposition of $B[p^{\infty}]\otimes k$ into a product ${\mathscr{G}}_{1}\times{\mathscr{G}}_{2}$. By Dieudonné-Manin classification ${\mathscr{G}}_{1}\times{\mathscr{G}}_{2}$ is isogenous to ${\mathscr{G}}_{1,g-1}\times{\mathscr{G}}_{g-1,1}$. Let $\psi$ be such an isogeny. Since the endomorphism ring preserves the kernel of $\psi$, it must also factor into a direct summand. This shows that ${\mathscr{G}}_{1}$ (respectively ${\mathscr{G}}_{2}$) is isogenous hence to ${\mathscr{G}}_{1,g-1}$ (respectively ${\mathscr{G}}_{g-1,1}$). By Lemma 4.4, we see that they’re in fact isomorphic. This shows that the isogeny $\psi$ is an isomorphism and hence $B$ lies in the central leaf of ${\mathscr{G}}_{1,g-1}\times{\mathscr{G}}_{g-1,1}$. ∎
@@ 5. Isogeny classes and the class group
Let $R\subset K$ be any order. Let $I\subset K$ be an invertible $R$ module. That is, $I$ is a fractional ideal of $R_{n}$ such that there exists some fractional ideal $J$ with $IJ=R$. We say two fractional ideals, $I_{1}$ and $I_{2}$ are equivalent if and only if there exists some $a\in K^{\times}$ such that $I_{1}=aI_{2}.$
@@
(Class group of an order) Under the equivalence mentioned above, we define the equivalence classes of fractional ideals of $R$ to be the class group of $R$, denoted as $Cl(R).$
The heart of rest of the paper lies in the following proposition that gives us a convenient characterisation of (isogenous) abelian varieties over $L^{\prime}$ with endomorphism ring $R_{n}$ in Definition 3.3. Here we recall the setup and notation at the beginning of Section 2. In particular, we return to the set up in characteristic zero with $\widetilde{A}$ an abelian variety over $L=W({\mathbb{F}}_{q})(1/p)$. Let $\Phi$ be the primitive CM type on $K$ corresponding $\widetilde{A}$ over $L$. We denote by $A$ over ${\mathbb{F}}_{q}$ its reduction mod $p$.
@@
For any positive integer $n$, there exists an injective map from the class group of $R_{n}$ to the set of isomorphism classes of abelian varieties over $L^{\prime}$ with endomorphism ring $R_{n}\subset(K,\Phi).$
@@
Let $I$ be an element of $Cl(R_{n}).$ We recall that $\Phi$ gives is an isomorphism $K\otimes{\mathbb{C}}={\mathbb{C}}^{\Phi}$ where ${\mathbb{C}}^{\Phi}$ denotes $g$ copies of ${\mathbb{C}}$ indexed by the embeddings in $\Phi.$ Since $I$ is a lattice in $K,$ we can embed it inside ${\mathbb{C}}^{\Phi}$. We refer to this embedding as $\Phi(I).$ The complex torus ${\mathbb{C}}^{\Phi}/\Phi(I)$ admits a polarisation (see Section 7) and therefore we get an abelian variety over ${\mathbb{C}}$ which we denote by $A_{\Phi(I)}$. Since $\Phi$ is primitive, $A_{\Phi(I)}$ is simple and so its endomorphism algebra is $K$. By construction $\operatorname{End}(A_{\Phi(I)})=R_{n}.$ Furthermore, since $A_{\Phi(I)}$ is a CM abelian variety, it descends, with its entire endomorphism algebra to a number field. Now, if the abelian variety had CM by the maximal order, then the abelian variety is defined over the maximal unramified extension of the reflex field $K^{*}$ and hence it would be defined over an unramified extension of ${\mathbb{Q}}_{p}.$ In our case, however, the field of definition might be slightly larger than that and may be a ramified extension of ${\mathbb{Q}}_{p}$. We first rule out this possibility for the abelian variety $A_{\Phi(R_{n})}.$ We note that $A_{\Phi(R_{n})}$ is isogenous to the abelian variety $A^{\prime}$ corresponding to the lattice ${\mathcal{O}}_{K}$ in $K$. Now, $A^{\prime}$ has CM by a maximal order over ¯$\overline{{\mathbb{Q}}}_{p}.$ Let $F$ be the unramified extension of ${\mathbb{Q}}_{p}$ where $A^{\prime}$ is defined. Since the kernel of the isogeny has order equal to the index of $R_{n}$ in ${\mathcal{O}}_{K}$, we can embed it inside $A^{\prime}[m]$ for some $(m,p)=1$. Now the kernel is defined over some extension $F^{\prime}$ of $F$. This must be unramified because the inertia group $I(\overline{v})$ acts trivially on the $m$-torsion and therefore on the subgroup as well. Here $\bar{v}$ refers to the extension of a place $v$ of $F^{\prime}$ to its separable closure. This descends $A_{\Phi(R_{n})}$ to an unramified extension of ${\mathbb{Q}}_{p}.$ This same argument works in descending an abelian variety corresponding to a fractional ideal $I$ whose index in $R_{n}$ is prime to $p.$ So suppose that $P$ is a fractional ideal of $R_{n}$ with index a power of $p$. We extend this to a fractional ideal $P^{\prime}$ of ${\mathcal{O}}_{K}$. Since $R_{n}$ is maximal at $p$, the index of $P$ in $R_{n}$ equals the index of $P^{\prime}$ in ${\mathcal{O}}_{K}$. Hence the index of $P$ in $P^{\prime}$ is prime to $p.$ Since the abelian variety corresponding to $P^{\prime}$ has endomorphism by the maximal order, it descends to an unramified extension of ${\mathbb{Q}}_{p}.$ We can now repeat above argument with $A_{\Phi(P)}$. We note that all the abelian varieties constructed above are isogenous to $\widetilde{A}_{L^{\prime}}$ over $L^{\prime}$ by Lemma 2.1. Finally, under the equivalence of ideals in $Cl(R_{n})$, two equivalent ideals would correspond to isomorphic lattices and hence give us isomorphic abelian varieties. ∎
Corollary 5.3.
Let $n$ be any positive integer. Then there exists an injective map from the class group $Cl(R_{n})$ to the set of isomorphism classes of abelian varieties over ${\mathbb{F}}_{q^{n}}$ isogenous to $A$ (without any condition on polarisation).
Proof.
@@
By Proposition 2.7, the reduction mod $p$ map from the set of isomorphism classes of abelian varieties isogenous to $\widetilde{A}$ over $L^{\prime}$ to that of $A$ over $k$ is injective. Moreover, for any $n$, by Lemma 4.1 the abelian variety constructed using an ideal $I\subset R_{n}$ is defined over $W({\mathbb{F}}_{q^{n}})(1/p)$ and hence the reduction mod $p$ of their Néron models descend to ${\mathbb{F}}_{q^{n}}$. By Lemma 2.2, their endomorphism ring is $R_{n}$ and hence have $\alpha^{n}\in K$ as their Frobenius morphism. ∎